# Randomness, Not Representation: The Unreliability of Evaluating Cultural Alignment in LLMs

### v1

@@ Abstract.
Research on the ‘cultural alignment’ of Large Language Models (LLMs) has emerged in response to growing interest in understanding representation across diverse stakeholders. Current approaches to evaluating cultural alignment borrow social science methodologies but often overlook systematic robustness checks. Here, we identify and test three assumptions behind current evaluation methods: (1) Stability: that cultural alignment is a property of LLMs rather than an artifact of evaluation design, (2) Extrapolability: that alignment with one culture on a narrow set of issues predicts alignment with that culture on others, and (3) Steerability: that LLMs can be reliably prompted to represent specific cultural perspectives. Through experiments examining both explicit and implicit preferences of leading LLMs, we find a high level of instability across presentation formats, incoherence between evaluated versus held-out cultural dimensions, and erratic behavior under prompt steering. We show that these inconsistencies can cause the results of an evaluation to be very sensitive to minor variations in methodology. Finally, we demonstrate in a case study on evaluation design that narrow experiments and a selective assessment of evidence can be used to paint an incomplete picture of LLMs’ cultural alignment properties. Overall, these results highlight significant limitations of current approaches for evaluating the cultural alignment of LLMs.
@@
Figure 1. Core assumptions about LLM cultural alignment fail under systematic evaluation. Our experiments reveal that cultural alignment in LLMs is: (1) not stable (Section 5.1 , Figure 2, Figure 3, Figure 4, Figure 5, Figure 6) - response variations from trivial format changes often exceed real-world cultural differences; (2) not extrapolable (Section 5.2 , Figure 7) - extrapolation from limited dimensions produces near-random clustering results, with strong sensitivity to which dimensions are included; and (3) not steerable (Section 5.3 , Figure 8) - even optimized prompting techniques produce erratic, un-humanlike response patterns that fail to align with cultural perspectives.
@@ 1. Introduction
Despite advancements in aligning AI with human preferences through methods like RLHF (Kovač et al., 2023; Casper et al., 2023), evaluating and controlling AI systems’ alignment with different cultures remains challenging (Prabhakaran et al., 2022; Ożegalska-Łukasik and Łukasik, 2023). Most existing research on evaluating cultural alignment in LLMs involves simply analyzing their responses to questions about their preferences (Pawar et al., 2024). For example, the CDEval benchmark evaluates cultural dimensions in LLMs using questions designed to assess six cultural dimensions (Wang et al., 2024b). However, modern LLMs – and their expressed ‘preferences’ – are complex (Liu et al., 2023). This leads us to ask whether current evaluation methods can adequately characterize the cultural alignment of LLMs.
Current methods for assessing cultural alignment borrow frameworks from social science methodologies (Lindgren and Holmström, 2020) but typically reduce to asking LLMs questions about their ‘preferences’ without necessarily implementing the systematic validation practices that establish robust findings. This oversight is potentially concerning because LLMs, unlike humans whose values often show consistent patterns, are known to express inconsistent ideas across different contexts (Sclar et al., 2023; Zheng et al., 2023; Errica et al., 2024; Zhuo et al., 2024; Anagnostidis and Bulian, 2024). Without proper validation, conclusions about LLMs’ cultural alignment may reflect evaluation design choices rather than the models’ actual properties.
Here, we investigate the reliability of current survey-based cultural alignment evaluations by identifying and testing three key assumptions from prior works (see Figure 1):
@@
  • Extrapolability: Alignment with one culture on a narrow set of issues predicts alignment with that culture on other issues.
@@
    We observe three assumptions in the cultural alignment evaluation literature: stability, extrapolability, and steerability.
@@
    We reveal significant inconsistencies in these widely held assumptions through a series of experiments involving explicit and implicit preference elicitation.
@@
    We demonstrate how subtle variations in evaluation methodology can dramatically shift conclusions about LLMs’ cultural alignment. We also present a case study showing that high-level findings on LLM preferences by Mazeika et al. (2025) do not replicate when the LLM is evaluated with the option of expressing indifference between alternatives.
Our findings suggest that current, popular methods for evaluating cultural alignment in LLMs require critical re-examination, as they risk oversimplifying or misrepresenting results. The high sensitivity of LLM responses to subtle methodological choices indicates that narrow experiments or a selective assessment of evidence may paint an incomplete picture of these systems’ cultural alignment properties. These results point to opportunities for developing more robust evaluation frameworks that better capture the complex and variable nature of LLM behavior.
@@ 2.1. Cultural Alignment in Language Models
An LLM’s ‘alignment’ with a culture describes the degree to which its behaviors reflect common beliefs within that culture regarding what is considered desirable and proper (Pawar et al., 2024). As language models have become increasingly capable of generating human-like text, researchers have focused on analyzing how the behaviors of these models reflect or diverge from different cultural perspectives (Cao et al., 2023). Some early work in cultural alignment has highlighted that LLMs tend to reflect values aligned with WEIRD (Western, Educated, Industrialized, Rich, and Democratic) societies (Johnson et al., 2022), leading to the perpetuation of cultural biases and the misrepresentation of non-WEIRD worldviews. This recognition has led researchers to adapt social science methodologies to evaluate cultural representation in LLMs (Guo et al., 2023).
@@ 2.2. Evaluating Cultural Alignment
Two main methodological approaches have emerged for assessing cultural alignment. The more common paradigm has been discriminative assessment, which adapts traditional survey techniques by requiring models to select from predetermined options to evaluate their preferences and biases (AlKhamissi et al., 2024; Arora et al., 2023; Cao et al., 2023; Masoud et al., 2023; Moore et al., 2024). For example, recent work by Mazeika et al. (2025) studies LLM ‘preferences’ by having them select a binary preference between two outcomes (e.g., “saving 10 lives in the United States” versus “saving 10 lives in Nigeria”).
In contrast, generative approaches analyze free-form model outputs, similar to qualitative methods in social science research (Adilazuarda et al., 2024; Tao et al., 2024). These can be implemented through both single-turn assessment, where cultural context and probe are given in one prompt, and multi-turn assessment, which evaluates responses over several interactions (Adilazuarda et al., 2024).
@@ 2.3. Survey-Based Assessments
Building on discriminative assessment, researchers have developed benchmarks to evaluate model responses across cultural contexts, from politics to religion (Durmus et al., 2023). These include the Value Survey Module (VSM) (Hofstede and Minkov, 2013) which assesses six cultural dimensions through a structured 24-question survey, and the Global Opinion Q&A dataset (GQA) (Durmus et al., 2023) which combines World Values Survey (WVS) and Pew Research questions to assess model responses across cultural contexts (Durmus et al., 2023).
Based on these surveys, recent work has introduced benchmarks for assessing cultural alignment in LLMs (AlKhamissi et al., 2024; Arora et al., 2023; Cao et al., 2023; Masoud et al., 2023; Wang et al., 2024b; Zhao et al., 2024; Tao et al., 2024; Kovač et al., 2023). However, prior studies examining survey-based approaches have revealed important methodological limitations. For instance, Arora et al. (2023) found that while multilingual pre-trained language models can learn cross-cultural value differences, their study found a weak correlation with human survey responses. Additionally, researchers have highlighted how LLMs’ preferences can be sensitive to prompting (Röttger et al., 2024). This suggests that survey methods can fail to capture and contextualize the nuances of LLMs’ cultural preferences.
@@ 3. Identifying Key Assumptions
In this section, we identify three foundational assumptions that underlie current approaches to evaluating cultural alignment in LLMs.
@@ 3.1. Stability
Stability Assumption Cultural alignment manifests as a property of LLMs that remains consistent across semantic-preserving variations in evaluation methodology, rather than being primarily an artifact of specific prompt design choices.
This assumption does not claim LLM outputs never vary with prompt changes (which would be trivially false). Rather, it posits that superficial evaluation design choices should not dramatically alter conclusions about an LLM’s cultural values, just as survey methodology variations should not fundamentally change our understanding of human cultural values.
@@
However, emerging evidence prompts us to question whether cultural alignment is a stable property of an LLM, suggesting that LLM behaviors and preferences can be sensitive to minor variations in prompt design (Sclar et al., 2023; Zheng et al., 2023; Errica et al., 2024; Zhuo et al., 2024; Anagnostidis and Bulian, 2024; Ceron et al., 2024; Dominguez-Olmedo et al., 2024; Gupta et al., 2024; tse Huang et al., 2024). Wang et al. (2024b) systematically quantified cultural response variations across different prompt formats, finding that models exhibited varying degrees of stability across three question styles, leading them to use weighted averaging to account for template-specific instabilities. Röttger et al. (2024) revealed that small shifts in prompt phrasing can produce variations larger than the differences between distinct real-world cultural positions. Building on this prior work, which has shown basic prompt sensitivity using direct questions about values, in this paper, we systematically examine how different evaluation approaches, including assessment of implicit biases, affect cultural response patterns.
@@ 3.2. Extrapolability
Extrapolability Assumption Alignment with one culture on a narrow set of issues reliably predicts alignment with that culture on other unobserved issues, such that a limited sample of cultural dimensions is sufficient to characterize an LLM’s overall cultural alignment.
Previous research evaluating the cultural alignment of LLMs has frequently used cultural dimensional frameworks, such as Hofstede’s cultural dimensions, [^1] Power distance, individualism vs. collectivism, masculinity vs. femininity, uncertainty avoidance, long-term orientation vs. short-term orientation, and indulgence vs. restraint. to evaluate alignment. This offers a useful lens into LLM preferences, but studies (Arora et al., 2023; Cao et al., 2023; Masoud et al., 2023) have identified consistently weak correlations between LLM outputs and established cultural value surveys, while others (Benkler et al., 2023; Motoki et al., 2024) found contradictory patterns in alignment across different values. Based on apparent LLM alignment with certain cultures on specific issues, some past works have concluded that LLMs align with those cultures in general (Santurkar et al., 2023; Benkler et al., 2023; Masoud et al., 2023) despite evidence of inconsistent alignment patterns.
However, evidence from both human behavior and recent LLM evaluations casts doubt on extrapolability. In human studies, cultural dimensions exhibit only partial correlation, with substantial variation driven by unique country-specific factors (Beugelsdijk and Welzel, 2018). These findings indicate that cultures cannot be reliably classified or characterized using a small number of cultural dimensions. Language models can demonstrate similar patterns of unexpected preferences across cultural dimensions. For example, Tao et al. (2024) found that while GPT models consistently exhibit self-expression biases, they display considerable variation between secular versus traditional values. Meanwhile, Wang et al. (2024b) observed that GPT-4’s high uncertainty avoidance in the wellness domain contrasts with its general trend. Similarly, in the context of political and ideological alignment, Santurkar et al. (2023) found that typically ‘liberal’ models like text-davinci-002 and text-davinci-003 express notably ‘conservative’ views on religious topics. In this paper, we build on prior work to statistically investigate the interplay between different cultural dimensions and test whether alignment on certain dimensions reliably predicts alignment on others in both LLMs and humans.
@@ 3.3. Steerability
Steerability Assumption Sufficiently detailed and context-rich instructions can reliably prompt an LLM to adopt coherent cultural stances that accurately reflect specific human cultural perspectives, allowing consistent modulation of an LLM’s cultural expression through prompting.
Prompting remains the most common method for steering LLMs toward specific behaviors or cultural perspectives, often referred to as “persona modulation” in the literature (Arora et al., 2023; Shah et al., 2023; Pawar et al., 2024; Niszczota et al., 2024; Jiang et al., 2023). For instance, AlKhamissi et al. (2024) introduced ‘anthropological prompting,’ which incorporates cultural context and reasoning frameworks to “shift responses toward cultural norms of underrepresented personas in Egyptian and American contexts.”
Despite past efforts, it is unclear the extent to which LLMs can reliably embody the assigned persona. One cause for doubt is that LLM fine-tuning methods (e.g., RLHF) tend to optimize for annotator approval, which is not guaranteed to make LLMs take on accurate personas (Durmus et al., 2023; Casper et al., 2023). Meanwhile, Tao et al. (2024) observe that cultural prompting can increase alignment for some countries while failing or even exacerbating bias for others. Meanwhile, Kovač et al. (2023) demonstrate that different “perspective induction” techniques can yield inconsistent results across tasks and model architectures. In this paper, we go beyond evaluating steerability in narrow contexts to show fundamental failures of LLMs to express humanlike preferences, let alone embody specific cultural perspectives, even under prompt optimization.
@@ 4. Experimental Setup
In this section, we present our experimental design for testing the three core assumptions of current cultural alignment evaluations in LLMs: stability, extrapolability, and steerability.
4.1. Model Selection and Configuration
We evaluated cultural alignment across five state-of-the-art LLMs: GPT-4o, Claude 3.5 Sonnet (20241022), Gemini 2.0 Flash, Llama 3.1 405B, and Mistral Large 2411. These models represent diverse architectures, releases (closed- and open-weight), and developers. To ensure deterministic outputs, all experiments used zero-temperature settings.
@@ 4.2. Cultural Alignment Assessment Frameworks
We evaluated LLM cultural alignment using various question/answer surveys as described below. Specific prompts are in Appendix A. All tests involved asking LLMs questions and eliciting responses on a Likert scale (e.g., strongly disagree, slightly disagree, neutral, slightly agree, strongly agree). For all numerical analysis, we continuously embedded Likert-scale responses using evenly-spaced values from 0 to 1 (e.g., 0.0, 0.25, 0.5, 0.75, 1.0).

### v2

@@ Abstract.
Research on the ‘cultural alignment’ of Large Language Models (LLMs) has emerged in response to growing interest in understanding representation across diverse stakeholders. Current approaches to evaluating cultural alignment through survey-based assessments that borrow from social science methodologies often overlook systematic robustness checks. Here, we identify and test three assumptions behind current survey-based evaluation methods: (1) Stability: that cultural alignment is a property of LLMs rather than an artifact of evaluation design, (2) Extrapolability: that alignment with one culture on a narrow set of issues predicts alignment with that culture on others, and (3) Steerability: that LLMs can be reliably prompted to represent specific cultural perspectives. Through experiments examining both explicit and implicit preferences of leading LLMs, we find a high level of instability across presentation formats, incoherence between evaluated versus held-out cultural dimensions, and erratic behavior under prompt steering. We show that these inconsistencies can cause the results of an evaluation to be very sensitive to minor variations in methodology. Finally, we demonstrate in a case study on evaluation design that narrow experiments and a selective assessment of evidence can be used to paint an incomplete picture of LLMs’ cultural alignment properties. Overall, these results highlight significant limitations of current survey-based approaches to evaluating the cultural alignment of LLMs and highlight a need for systematic robustness checks and red-teaming for evaluation results. Data and code are available at akhan02/cultural-dimension-cover-letters and ariba-k/llm-cultural-alignment-evaluation, respectively.
@@
Figure 1. Core assumptions about LLM cultural alignment fail under systematic evaluation. Our experiments reveal that cultural alignment in LLMs is: (1) not stable (Section 5.1, Figure 3, Figure 4) - response variations from trivial format changes often exceed real-world cultural differences; (2) not extrapolable (Section 5.2, Figure 5) - extrapolation from limited dimensions produces near-random clustering results, with strong sensitivity to which dimensions are included; and (3) not steerable (Section 5.3, Figure 6) - even optimized prompting techniques produce erratic, un-humanlike response patterns that fail to align with cultural perspectives.
@@ 1. Introduction
Despite advancements in aligning AI with human preferences through methods like RLHF (Kovač et al., 2023; Casper et al., 2023), evaluating and controlling AI systems’ alignment with different cultures remains challenging (Prabhakaran et al., 2022; Ożegalska-Łukasik and Łukasik, 2023). Most existing research evaluating cultural alignment in LLMs uses survey-based assessments that merely analyze how models respond when asked about their values and preferences (Pawar et al., 2024). For example, the CDEval benchmark evaluates cultural dimensions in LLMs using questions designed to assess six cultural dimensions (Wang et al., 2024b). However, modern LLMs – and their expressed ‘preferences’ – are complex (Liu et al., 2023). This leads us to ask whether current survey-based evaluation methods can adequately characterize the cultural alignment of LLMs.
Current methods for assessing cultural alignment borrow frameworks from social science methodologies (Lindgren and Holmström, 2020) but typically reduce to asking LLMs questions about their ‘preferences’ without necessarily implementing systematic validation practices. This oversight is potentially concerning because LLMs, unlike humans whose values often show consistent patterns, are known to express inconsistent ideas across different contexts (Sclar et al., 2023; Zheng et al., 2023; Errica et al., 2024; Zhuo et al., 2024; Anagnostidis and Bulian, 2024).
Here, we investigate the reliability of current survey-based cultural alignment evaluations by identifying and testing three key assumptions from prior work (see Figure 1):
@@
  • Extrapolability: Alignment with one culture on a narrow set of issues predicts alignment with it on others.
@@
    We identify and formally characterize three critical assumptions in survey-based cultural alignment evaluation literature: stability, extrapolability, and steerability.
@@
    We provide comprehensive empirical evidence challenging these assumptions through experiments with both explicit surveys and implicit preference elicitation, quantifying significant inconsistencies across different cultural dimensions.
@@
    We demonstrate how subtle variations in evaluation methodology can greatly shift conclusions about LLMs’ cultural alignment, presenting a case study that demonstrates how high-level findings from Mazeika et al. (2025) do not replicate when an LLM is evaluated with the option of selecting indifference between alternatives.
Our findings suggest that current, popular survey-based methods for evaluating cultural alignment in LLMs require critical re-examination, as they risk oversimplifying or misrepresenting results. The high sensitivity of LLM responses to subtle methodological choices indicates that narrow experiments or a selective assessment of evidence may paint an incomplete picture of these systems’ cultural alignment properties. We release data and code at akhan02/cultural-dimension-cover-letters and ariba-k/llm-cultural-alignment-evaluation, respectively.
@@ 2.1. Cultural Alignment in Language Models
An LLM’s ‘alignment’ with a culture describes the degree to which its behaviors reflect common beliefs within that culture regarding what is considered desirable and proper (Pawar et al., 2024). As language models have become increasingly capable of generating human-like text, researchers have analyzed how the behaviors of these models reflect or diverge from different cultural perspectives (Cao et al., 2023). Early work in cultural alignment has highlighted that LLMs tend to reflect values aligned with WEIRD (Western, Educated, Industrialized, Rich, and Democratic) societies (Johnson et al., 2022), leading to the perpetuation of cultural biases and the misrepresentation of non-WEIRD worldviews.
@@ 2.2. Evaluating Cultural Alignment
Two main methodological approaches have emerged for assessing cultural alignment. The more common paradigm has been discriminative assessment, which adapts traditional survey techniques by requiring models to select from predetermined options to evaluate their preferences and biases (AlKhamissi et al., 2024; Arora et al., 2023; Cao et al., 2023; Masoud et al., 2023; Moore et al., 2024). For example, recent work by Mazeika et al. (2025) studies LLM ‘preferences’ by having them select a binary preference between two outcomes (e.g., “saving 10 lives in the United States” versus “saving 10 lives in Nigeria”). In contrast, generative approaches analyze free-form model outputs, similar to qualitative methods in social science research (Adilazuarda et al., 2024; Tao et al., 2024). These can be implemented through both single-turn assessment, where cultural context and probe are given in one prompt, and multi-turn assessment, which evaluates responses over several interactions (Adilazuarda et al., 2024).
@@ 2.3. Survey-Based Assessments
Building on discriminative assessment, researchers have developed benchmarks to evaluate model responses across cultural contexts, from politics to religion (Durmus et al., 2023). These include the Value Survey Module (VSM) (Hofstede and Minkov, 2013) which assesses six cultural dimensions through a structured 24-question survey, and the Global Opinion Q&A dataset (GQA) (Durmus et al., 2023) which combines World Values Survey (WVS) and Pew Research questions to assess views across cultural contexts (Durmus et al., 2023).
Based on these surveys, recent work has introduced benchmarks for assessing cultural alignment in LLMs (AlKhamissi et al., 2024; Arora et al., 2023; Cao et al., 2023; Masoud et al., 2023; Wang et al., 2024b; Zhao et al., 2024; Tao et al., 2024; Kovač et al., 2023). However, prior studies on survey-based approaches have shown methodological limitations. For instance, Arora et al. (2023) found that while LLMs can reflect cross-cultural value differences, there was a weak correlation with human survey responses. Other work has highlighted how LLMs’ preferences can be sensitive to prompting (Röttger et al., 2024). This suggests that survey methods may struggle to capture and contextualize the nuances of LLMs’ cultural preferences.
@@ 3. Identifying Key Assumptions
Here, we identify three assumptions that underlie current approaches to evaluating cultural alignment in LLMs.
@@ 3.1. Stability
Stability Assumption Cultural alignment manifests as a property of LLMs that generally remains consistent across semantic-preserving variations in evaluation methodology, rather than being primarily an artifact of specific prompt design choices.
@@
However, emerging evidence suggests that cultural alignment may not be a stable property of an LLM, as behaviors and preferences can be sensitive to minor variations in prompt design (Sclar et al., 2023; Zheng et al., 2023; Errica et al., 2024; Zhuo et al., 2024; Anagnostidis and Bulian, 2024; Ceron et al., 2024; Dominguez-Olmedo et al., 2024; Gupta et al., 2024; tse Huang et al., 2024). Wang et al. (2024b) quantified cultural response variations across different prompt formats, finding that models exhibited varying degrees of stability across question styles. Röttger et al. (2024) revealed that small shifts in prompt phrasing can produce variations larger than the differences between preferences. Building on this prior work that has demonstrated basic prompt sensitivity using direct questions about values, in this paper we examine how different evaluation approaches, including assessment of implicit biases, affect measurements of cultural alignment.
@@ 3.2. Extrapolability
Extrapolability Assumption Alignment with one culture on a narrow set of issues generally predicts alignment with that culture on other unobserved issues, such that a limited sample of cultural dimensions is sufficient to characterize an LLM’s overall cultural alignment.
Previous research evaluating the cultural alignment of LLMs has frequently used cultural dimensional frameworks, such as Hofstede’s cultural dimensions, [^1] These dimensions include: power distance index, individualism vs. collectivism, masculinity vs. femininity, uncertainty avoidance index, long-term vs. short-term orientation, and indulgence vs. restraint. to evaluate alignment. This offers a useful lens into LLM preferences, but studies (Arora et al., 2023; Cao et al., 2023; Masoud et al., 2023) have identified consistently weak correlations between LLM outputs and established cultural value surveys. Nevertheless, several studies have suggested broader cultural alignment of LLMs based on observations of alignment on specific cultural issues (Santurkar et al., 2023; Benkler et al., 2023; Masoud et al., 2023).
However, evidence from both human behavior and recent LLM evaluations calls extrapolability into question. In human studies, cultural dimensions exhibit only partial correlation, with substantial variation driven by unique country-specific factors (Beugelsdijk and Welzel, 2018). Meanwhile, language models can demonstrate similar patterns of unexpected preferences across cultural dimensions. For example, Tao et al. (2024) found that while GPT models consistently exhibit self-expression biases, they also display considerable variation between secular versus traditional values. Similarly, in the context of political and ideological alignment, Santurkar et al. (2023) found that typically ‘liberal’ models like text-davinci-002 and text-davinci-003 express notably ‘conservative’ views on religious topics. In this paper, we build on prior work to statistically investigate the interplay between different cultural dimensions and test whether alignment on certain dimensions reliably predicts alignment on others in both LLMs and humans.
@@ 3.3. Steerability
Steerability Assumption LLMs can be reliably prompted to embody coherent cultural stances that accurately reflect specific human cultural perspectives.
Prompting remains the most common method for steering LLMs toward specific behaviors or cultural perspectives, often referred to as “persona modulation” in the literature (Arora et al., 2023; Shah et al., 2023; Pawar et al., 2024; Niszczota et al., 2024; Jiang et al., 2023). For example, AlKhamissi et al. (2024) introduced ‘anthropological prompting, which incorporates cultural context and reasoning frameworks to “shift responses toward cultural norms of underrepresented personas in Egyptian and American contexts.”
Despite past efforts, it is unclear to what extent LLMs can reliably embody the assigned persona. One reason for doubt is that LLM fine-tuning methods (e.g., RLHF) tend to optimize for annotator approval, which is not guaranteed to make LLMs take on accurate personas (Durmus et al., 2023; Casper et al., 2023). Tao et al. (2024) observe that cultural prompting can increase alignment for some countries while failing or even exacerbating bias for others. Additionally, Kovač et al. (2023) show that different “perspective induction” techniques can yield inconsistent results across tasks and model architectures. In this paper, we go beyond evaluating steerability in narrow contexts to show fundamental failures of LLMs to express humanlike preferences, let alone embody specific cultural perspectives, even under prompt optimization.
@@ 4. Experimental Setup
4.1. Model Selection, Temperature, and Experiment Configuration
For all experiments in Section 5, we evaluated cultural alignment across five state-of-the-art LLMs: GPT-4o (OpenAI), Claude 3.5 Sonnet 20241022 (Anthropic), Gemini 2.0 Flash (Google), Llama 3.1 405B (Meta), and Mistral Large 2411 (Mistral AI). These models represent diverse architectures, releases (closed- and open-weight), and developers. To ensure consistent and reproducible results, we applied the same experimental protocol: temperature was set to 0.0 for deterministic outputs, each query was run three independent times, and measurements were averaged across these trials. Any exceptions to this standard protocol are explicitly noted in the relevant subsections.
@@ 4.2. Cultural Alignment Assessment Frameworks
To evaluate the three key assumptions, we used the question/answer surveys described below. All experiments involved asking LLMs questions and eliciting responses on a Likert scale (e.g., strongly disagree, slightly disagree, neutral, slightly agree, strongly agree). For the specific prompts used in each experiment, see Appendix A.